One third of a number decreased by 2 is the same as that number increased by 6. Find that number.

Answer by stanbon(75887)

One third of a number decreased by 2 is the same as that number increased by 6. Find that number.
-------------------
Equation:
(1/3)x -2 = x+6
(1/3)x = x+8
Subtract (1/3)x from both sides.
Subtract 8 from both sides.
(2/3)x = -8
(1/3)x = -4
x = -12
===================
Cheers,
Stan H.